OpenStudy (anonymous):

quadratic form (polynomials) 15x^5-72x^3-108x help please?? medals given!

OpenStudy (anonymous):

oh my loard

OpenStudy (anonymous):

okay first i believe you can factor that all by 3

OpenStudy (anonymous):

Simplifying 15x5 + -72x3 + -108x = 0 Reorder the terms: -108x + -72x3 + 15x5 = 0 Solving -108x + -72x3 + 15x5 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '3x'. 3x(-36 + -24x2 + 5x4) = 0 Factor a trinomial. 3x((-6 + -5x2)(6 + -1x2)) = 0 Ignore the factor 3.

OpenStudy (anonymous):

your factored polynomial is 3x(x2−6)(5x2+6) when its all said and done

OpenStudy (anonymous):

oh god thanks so much!!!!!!